a bus travel at a distance of 90 km in 2 hours and another 30 km in 1 hour. what is the average speed of the bus?

Answer by Alan3354(69443)

a bus travel at a distance of 90 km in 2 hours and another 30 km in 1 hour. what is the average speed of the bus?
-------------
d = 90 + 30 = 120 km
t = 2 + 1 = 3 hours
---------
r = d/t